1. An infrared radiometer on a satellite measures radiation emitted from Earths surface in the atmospheric window part of
Answer #1

By Planck's law, spectral radiance per unit wavelength λ is given by

{\displaystyle B_{\lambda }(\lambda ,T)={\frac {2hc^{2}}{\lambda ^{5}}}{\frac {1}{e^{\frac {hc}{\lambda k_{\mathrm {B} }T}}-1}},}

where kB(Boltzmann constant) = 1.38×10−23 J/K ,

h (Planck constant)= 6.626×10−34 J.s

and c(speed of light in the medium) = 3 x 10 8 m/s

S.I. is of Bλ are W·sr−1·m−3

Bλ = 8 W·sr−1·m−2\mum-1 = 8 x 106  W·sr−1·m−3

and λ = 10 \mum = 10 x 10-6 m = 10-5 m

So , By above formula

| ره) ( G2 ( x64 یا , روم، 5 . 38 اع x 2ر x 3x1 رو3-8x10 3x . 3x16 ا گر 5-10) 1 وا x 9 . 264 ا ا ب ا ۲ 4 - م اہا. ۱۹۹۵ به . x

Average Earth's temperature is about 287.  Air temperature drops 9.84 K per 1000 m of altitude. Hence we can say that Cloud height should be around 1 Km.